# 20xx Vol. xx No. x, XXX

### v1

@@ 3.2 Strain energy during global quakes
The gravitational and kinetic energy changes correspond to the shape change and spin-down. However, the ellipticity change during the glitch is small, and the total energy for the glitch epoch is approximated as that of the glitch phase. The total energy change during the quake is mainly the strain energy, which is calculated as (Zhou et al. 2014)
@@
The upper limit for the total strain energy is the total spin energy (see Section 3.3).
@@
The waiting time for the starquakes is much longer than the waiting time for some short-interval FRBs of $\sim 10\,\rm ms-100\,s$. Strain energy is therefore not a viable FRB power source since it could not release frequently.
3.3 Rotational energy of spindown
The rotational energy of a neutron star can sustain radio emission from a pulsar (Gold 1968). The strong magnetic field and high rotation speed of neutron stars cause the magnetosphere to fill with a plasma, leading to radiation in the pattern of a rotating beacon (Goldreich & Julian 1969). Regardless of whether the radiation mechanisms of pulsars and FRBs are the same, we consider the possibility that both of them are powered by rotation. The total spin energy of a rotating neutron star with $M=1.4M_{\odot}$ is given by
@@ 3.3 Rotational energy of spindown
where we do not consider an inclination angle dependence because the power is insensitive to the angle when both magnetic dipolar spindown and wind spindown are considered (Harding et al. 1999; Xu & Qiao 2001; Spitkovsky 2006) The corresponding radiation luminosity should be equal to the spin-down power, which gives the period evolution of the neutron star:
@@
The spin-down age is $\approx P/(2\dot{P})$ when the initial rotational period $P_{0}$ is much smaller than $P$.
For a typical magnetar with $P\sim 1$ s, the rotational energy is in the same order as the total magnetic energy stored in the crust, given as Equation (7). According to discussions in Section 3.1, the rotational energy can sustain $\sim 10^{4}$ bursts. However, the spin-down power of a magnetar with $P\sim 1$ s may not be able to support very bright bursts whose observed isotropic luminosities greatly exceed the spindown luminosity.
We consider that a magnetar with $P=0.1$ s, which is the point of interest in the following discussion. The spin-down power is sufficient for the FRB luminosity. However, the spin-down age is calculated to be $\sim$ 0.2 years which is much smaller than the age since the FRBs were discovered. There might be a lot of bursts missing before the FRB source was first discovered. The spindown age of several years indicates that the current period of the magnetic star is $\sim 0.5$ s. As the rapid spindown evolution, the luminosity of FRB decreases significantly.
Such a newly born magnetar is surrounded by a supernova remnant. The free-free optical depth for the wave traveling through the ejecta shell ($Z=8$, $A=16$) for an oxygen-dominated composition) is given by (Metzger et al. 2017)
@@
Figure 1: $P-B_{s}$ diagram. We take $M_{\rm ej}=10M_{\odot}$, $v_{\rm ej}=10^{4}\,\rm km\,s^{-1}$, $\nu=10^{9}$ GHz, $T_{\rm ej}=10^{4}$ K. The black dashed lines denote spin-down energy rates which $\dot{E}_{d}=10^{36},10^{38},10^{40}\,\rm erg\,s^{-1}$. The black solid lines denote spin-down ages which are 0.1, 10, and $10^{3}$ yr. The red and blue dashed-dotted lines denote $\tau_{C}=1$ and $\tau_{\rm ff}=1$. The red and blue regions denote the optically thick for induced Compton scattering and free-free absorption.
@@ 3.4 Gravitational and phase transition energies
where $\delta R$ is a global reduction in radius. When converting to hybrid stars, the phase transition and the associated energy release have been shown to be able to source core-quakes (Bejger et al. 2005) and repeated fast radio bursts (Shen et al. 2023). Recently, it was shown that a phase transition from strangeon matter  [^1] The building blocks of a strangeon star are strangeons which are localized quark clusters formed by strong force, in analogy of atomic nucleons but with a large baryon number and with strange quarks in almost equal fraction as up and down quarks. to strange quark matter is feasible in the core region of strangeon stars (Zhang et al. 2023), which may also result in core quakes and release of a large amount of energy due to the gravitational mass changes.
@@
The radiation mechanism from gravitational energy to radiation is not very determined. One possible mechanism is that the star has oscillations induced by seismic waves. The oscillations can enhance the voltage potential on the polar gap causing more energetic emissions (Lin et al. 2015). A minority of the released gravitational energy is transformed into kinetic energy of the oscillations and then becomes radiation power. However, it is not clear how much energy is converted into thermal energy during core heating via quakes. The gravitational energy budget seems to give a theoretical upper limit on the energy released from a pulsar-like compact star.
Another interesting scenario is relevant to the challenging equation of state of cold supra-nuclear matter, in which huge gravitational energy may efficiently power emission if the equation of state is so stiff that the maximum mass, $M_{\rm TOV}$, calculated with the Tolman-Oppenheimer-Volkoff equation reaches $(2.5\sim 3.0)M_{\odot}$, particularly after binary neutron star merger. The mass of a merger remnant could even be slightly higher than $M_{\rm TOV}$ for two companion stars with typical masses of $\sim 1.5M_{\odot}$, but the hyper-massive neutron star would survive due to rapid rotation before collapsing into a black hole. Besides the centrifugal force, in a strangeon star model, an elasticity would additionally prevent the remnant from collapsing after solidification (Yuan et al. 2017), and quakes could then occur frequently afterward. This kind of massive remnants could also be attractive gravitational wave sources, even echoes (Zhang et al. 2023a), with multi-messengers. In any cases, there are quantitative uncertainties in this scenario, to be investigated in the future.
Alternatively, another way to release gravitational energy is for neutron stars to accrete external matter, e.g., asteroid (Dai et al. 2016; Geng et al. 2021). The accreted matter falls onto the stellar surface leading to starquakes. During this process, the gravitational energy is firstly converted into kinetic energy and then transformed into emitting charges via somehow mechanisms. However, it is hard to interpret how frequently the bursts can happen, e.g., the event rate can be up to $100\,\rm h^{-1}$.
The energy evolutionary paths are summarized in Figure 2.
@@ 4.1 Narrow spectra and impact on energy distribution
Observations of repeaters show that the energy/luminosity function spans in a wide range. The distribution has a high energy cut-off, while no sharp cut-off at the low energy bands due to sensitivity limitations. For some radio telescopes, the limited sensitivity only allows to catch a small number of bright bursts whose function can be well modeled by a power law distribution. If the sensitivity is high enough, one can see hidden features.
@@
4.2 Period of FRBs
Many FRB models invoke rotating compact stars. However, the spin period has not been found from burst timing analyses (e.g., Niu et al. 2022b). FRBs are much more energetic than radio pulses from normal pulsars, so the well-defined open field line region is likely distorted and enlarged. Within the framework of a magnetar, a starquake can lead to twisting of magnetic field, and the twisted current-carrying bundle plays the role similar to the open field line region of normal pulsars. The observed duty cycles are then enlarged and the emission phase is more random, matching the observation of FRB-like bursts from SGR J1935+2154 (Zhu et al. 2023).
The missing period of FRBs may have a strong relationship with the trigger mechanism. We speculate that the charged emitting particles can be triggered by point discharges from “hills” on the stellar surface. The hills may be created by crust shear and motion which may be randomly distributed on the surface. Some hills are higher and sharper, allowing them to generate more charges, emitting brighter pulses. These higher hills may stand more stably during frequent crustal activities, unlike the shorter hills that may disappear. As a result, some bright bursts can have stable phases, and timing measurements of them could display spin period modulation (e.g., Li et al. 2024). As shown in Figure 1, if the magnetar was born from a core-collapse supernova, the spin period cannot be much shorter than $\sim 1$ s, unless the bursts are too faint. The non-detection of periodicity from FRBs does not necessarily rule out a rotating compact star as the FRB source.
@@ 4.3 Duty Cycle
We simply discuss the duty cycle by considering constant burst rate during the lifetime. However, for most repeaters, we are not sure whether they are physically inactive or due to the observational sensitivity, and the constant burst rate in terms of time is a strong hypothesis. Some repeaters may have periodically active windows (e.g., Chime/Frb Collaboration et al. 2020; Rajwade et al. 2020). If the truly duty cycle is small and a high value is adopted as the average, we may overestimate the energy budget. A detailed burst rate evolution in time needs to be considered.
@@ 5 Summary
In this work, we discuss the energy budget of FRBs and investigate various energy sources and dissipation mechanisms related to starquakes. We have drawn the following conclusions:
@@
2. The strain energy released during the quake is much smaller than the magnetic field energy. For a neutron star with a solid crust and a stiff equation of state, the strain energy can accumulate during neutron star spin-down and is suddenly released during a glitch. However, the star needs to spend a long time to accumulate an enough strain energy to induce a starquake. This time is much longer than the observed waiting times of FRBs.
3. The total rotational energy of a normal magnetar with $P\sim 1$ s is comparable to the magnetic energy stored in the crust, while the spin-down power cannot support some bright FRBs. The rotational energy of a magnetar whose period is comparable to 0.1 s or shorter can match the energy and luminosity budget of FRBs. However, the GHz radio emission cannot escape from the ejecta shell region for an extremely young magnetar due to the high density. The spindown timescale is also too short compared with the observed life times of active repeaters.
4. For the neutron star scenario, quakes of a quark matter core with phase transition can release gravitational energy and phase transition energy. For the solid quark star (strangeon star) scenario, starquakes can happen frequently and may release a huge amount of gravitational energy if the equation of state is so stiff that the maximum mass can be as high as $(2.5\sim 3.0)M_{\odot}$, allowing a stable merger product from a binary merger. The total gravitational energy is high but it is not known whether such energy can be converted to emission with a large enough efficiency. This energy budget seems to give a theoretical upper limit of energy release of a pulsar-like compact star.
@@ Acknowledgements.
We are grateful to Shunshun Cao, Mingyu Ge, Kejia Lee, Bing Li, Ze-Nan Liu, Yudong Luo, Chenhui Niu, Yuan-Pei Yang, Xiaoping Zheng, Xia Zhou, and members of “Dejiang Find Girlfriend Assistant” Wechat group for helpful discussion. This work is supported by the National SKA Program of China (No. 2020SKA0120100) and the Strategic Priority Research Program of the CAS (No. XDB0550300). J.F.L. acknowledges support from the NSFC (Nos.11988101 and 11933004) and from the New Cornerstone Science Foundation through the New Cornerstone Investigator Program and the XPLORER PRIZE.

### v2

@@ 3.3.1 Spindown-induced glitch with starquake
The gravitational and kinetic energy changes correspond to the shape change and spindown. However, the ellipticity change during the glitch is small, and the total energy for the glitch epoch is approximated as that of the glitch phase. The total energy change during the quake is mainly the strain energy, which is calculated as (Zhou et al. 2014)
@@
The upper limit for the total strain energy is the total spin energy.
@@
The waiting time for the starquakes is much longer than the waiting time for some short-interval FRBs of $\sim 10\,\rm ms-100\,s$. Strain energy is therefore not a viable FRB power source since it cannot be released frequently.
3.3.2 Rotational energy for young mangetars
Regardless of whether the radiation mechanisms of pulsars and FRBs are the same, we consider the spindown power of a neutron star to sustain FRBs. The spindown power is independent of a starquake scenario and is just regarded as an energy mechanism in the following discussion.
The rotational energy of a neutron star can sustain radio emission from a pulsar (Gold 1968). The strong magnetic field and high rotation speed of neutron stars cause the magnetosphere to fill with plasma, leading to radiation in the pattern of a rotating beacon (Goldreich & Julian 1969). The total spin energy of a rotating neutron star with $M=1.4M_{\odot}$ is given by
@@ 3.3.2 Rotational energy for young mangetars
where we do not consider an inclination angle dependence because the power is insensitive to the angle when both magnetic dipolar spindown and wind spindown are considered (Harding et al. 1999; Xu & Qiao 2001; Spitkovsky 2006). The corresponding radiation luminosity should be equal to the spindown power, which gives the period evolution of the neutron star:
@@
The spindown age is
$t_{\rm sd}=\tau_{\mathrm{c}}\left[1-\left(\frac{P_{0}}{P}\right)^{2}\right],$ (20)
where $P_{0}$ is the initial rotational period. It can be reduced to $\approx P/(2\dot{P})$ when the initial rotational period $P_{0}$ is much smaller than $P$.
For a typical magnetar with $P\sim 1$ s, the rotational energy is in the same order as the total magnetic energy stored in the crust, given as Equation (8). According to discussions in Section 3.3, the rotational energy can sustain $\sim 10^{4}$ bursts. However, the spindown power of a magnetar with $P\sim 1$ s may not be able to support very bright bursts whose observed isotropic luminosities greatly exceed the spindown luminosity.
We consider a magnetar with $P=0.1$ s, which is the point of interest in the following discussion. The spindown power is sufficient for the FRB luminosity. However, the spindown age is calculated to be $\sim$ 0.2 years which is much smaller than the age since the FRBs were discovered. There might be a lot of bursts missing before the FRB source was first discovered. As the rapid spindown evolution, the luminosity of FRB decreases significantly.
Such a newly born magnetar is surrounded by a supernova remnant. We take the spindown age as the age of supernova remnant. The free-free optical depth for the wave travelling through the ejecta shell ($Z=8$, $A=16$) for an oxygen-dominated composition) is given by (Metzger et al. 2017)
@@
Figure 3: $P-B_{s}$ diagram. We take $M_{\rm ej}=10M_{\odot}$, $v_{\rm ej}=10^{4}\,\rm km\,s^{-1}$, $\nu=10^{9}$ GHz, $T_{\rm ej}=10^{4}$ K. The black dashed lines denote spindown energy rates which $\dot{E}_{d}=10^{36},10^{38},10^{40}\,\rm erg\,s^{-1}$. The black solid lines denote spindown ages which are 0.1, 10, and $10^{3}$ yr. The red and blue dashed-dotted lines denote $\tau_{C}=1$ and $\tau_{\rm ff}=1$. The red and blue regions denote the optically thick for induced Compton scattering and free-free absorption.
@@ 3.4 Gravitational and phase transition energies
where $\delta R$ is a global reduction in radius. When converting to hybrid stars, the phase transition and the associated energy release have been shown to be able to source core-quakes (Bejger et al. 2005) and repeated fast radio bursts (Shen et al. 2023). Recently, it was shown that a phase transition from strangeon matter  [^2] The building blocks of a strangeon star are strangeons which are localized quark clusters formed by the strong force, in analogy of atomic nucleons but with a large baryon number and with strange quarks in almost equal fraction as up and down quarks. to strange quark matter is feasible in the core region of strangeon stars (Zhang et al. 2023), which may also result in core quakes and release of a large amount of energy due to the gravitational mass changes.
@@
The radiation mechanism from gravitational energy to radiation is not very determined. One possible mechanism is that the star has oscillations induced by seismic waves. The oscillations can enhance the voltage potential on the polar gap causing more energetic emissions (Lin et al. 2015). A part of the released gravitational energy is transformed into kinetic energy of the oscillations and then becomes radiation power. However, it is not clear how much energy is converted into thermal energy during core heating via quakes. The gravitational energy budget seems to give a theoretical upper limit on the energy released from a pulsar-like compact star.
Another interesting scenario is relevant to the challenging equation of state of cold supra-nuclear matter, in which huge gravitational energy may efficiently power emission if the equation of state is so stiff that the maximum mass, $M_{\rm TOV}$, calculated with the Tolman-Oppenheimer-Volkoff equation reaches $(2.5\sim 3.0)M_{\odot}$, particularly after binary neutron star merger. The mass of a merger remnant could even be slightly higher than $M_{\rm TOV}$ for two companion stars with typical masses of $\sim 1.5M_{\odot}$, but the hyper-massive neutron star would survive due to rapid rotations before collapsing into a black hole. Besides the centrifugal force, in a strangeon star model, an elasticity would additionally prevent the remnant from collapsing after solidification (Yuan et al. 2017), and quakes could then occur frequently afterward. This kind of massive remnants could also be attractive gravitational wave sources, even echoes (Zhang et al. 2023a), with multi-messengers. In any case, there are quantitative uncertainties in this scenario to be investigated in the future.
Alternatively, another way to release gravitational energy is for neutron stars to accrete external matter, e.g., asteroid (Dai et al. 2016; Zhang et al. 2018b; Geng et al. 2021). The accreted matter falls onto the stellar surface leading to starquakes. During this process, the gravitational energy is firstly converted into kinetic energy and then transformed into emitting charges via somehow mechanisms. However, it is hard to interpret how frequently the bursts can happen, e.g., the event rate can be up to $100\,\rm h^{-1}$.
@@ 4.1 Narrow spectra and impact on energy distribution
The bandwidth limitations may affect the calculation of burst energy. Observations of repeaters show that the energy/luminosity function spans in a wide range. The distribution has a high energy cut-off, while no sharp cut-off at the low energy bands due to sensitivity limitations. For some radio telescopes, the limited sensitivity only allows to catch a small number of bright bursts whose function can be well modelled by a power law distribution. If the sensitivity is high enough, one can see hidden features.
@@ 4.2 Duty Cycle
We simply discuss the duty cycle by considering constant burst rate during the lifetime. However, for most repeaters, we are not sure whether they are physically inactive or due to the observational sensitivity, and the constant burst rate in terms of time is a strong hypothesis. Some repeaters may have periodically active windows (e.g., Chime/Frb Collaboration et al. 2020; Rajwade et al. 2020). If the true duty cycle is small and a high value is adopted as the average, we may overestimate the energy budget. A detailed burst rate evolution in time needs to be considered.
4.3 Period for energetic FRBs?
Many FRB models invoke rotating compact stars. However, the spin period has not been found from burst timing analyses (e.g., Niu et al. 2022b). FRBs are much more energetic than radio pulses from normal pulsars, so the well-defined open field line region is likely distorted and enlarged. Within the framework of a magnetar, a starquake can lead to twisting of the magnetic field, and the twisted current-carrying bundle plays a role similar to the open field line region of normal pulsars. The front of fireballs due to the X-ray bursts can push aside the pair plasma of closed field lines, forming a clean region to let radio waves escape (Ioka 2020). The observed duty cycles are then enlarged and the emission phase is more random, matching the observation of FRB-like bursts from SGR J1935+2154 (Zhu et al. 2023).
The missing period of FRBs may have a strong relationship with the trigger mechanism. We speculate that the charged emitting particles can be triggered by point discharges from “hills” on the stellar surface. The hills may be created by crust shear and motion which may be randomly distributed on the surface. Some hills are higher and sharper, allowing them to generate more charges, and emitting brighter pulses. These higher hills may stand more stably during frequent crustal activities, unlike the shorter hills that may disappear. As a result, some bright bursts can have stable phases, and timing measurements of them could display spin period modulation (e.g., Li et al. 2024). As shown in Figure 3, if the magnetar was born from a core-collapse supernova, the spin period cannot be much shorter than $\sim 1$ s, unless the bursts are too faint. The non-detection of periodicity from FRBs does not necessarily rule out a rotating compact star as the FRB source.
@@ 5 Summary
In this work, we discuss the energy budget of FRBs and investigate various energy sources and dissipation mechanisms related to starquakes. During the starquake, the strain energy can be directly released and possibly accompanied by the other three kinds of energy releasing and conversion. We have drawn the following conclusions:
@@
2. For a neutron star with a solid crust and a stiff equation of state, the strain energy can accumulate during the neutron star spindown and is suddenly released during a glitch. However, the star needs to spend a long time to accumulate an enough strain energy to induce a starquake. This time is much longer than the observed waiting times of FRBs.
3. The total rotational energy of a normal magnetar with $P\sim 1$ s is comparable to the magnetic energy stored in the crust, while the spindown power cannot support some bright FRBs. The rotational energy of a magnetar whose period is comparable to 0.1 s or shorter can match the energy and luminosity budget of FRBs. However, the GHz radio emission cannot escape from the ejecta shell region for an extremely young magnetar due to the high density. The spindown timescale is also too short compared with the observed lifetimes of active repeaters.
4. For the neutron star scenario, quakes of a quark matter core from phase transitions can release gravitational energy and phase transition energy. For the solid quark star (strangeon star) scenario, starquakes can happen frequently and may release a huge amount of gravitational energy if the equation of state is so stiff that the maximum mass can be as high as $(2.5\sim 3.0)M_{\odot}$, allowing a stable merger product from a binary merger. The total gravitational energy is high but it is not known whether such energy can be converted to emissions with a large enough efficiency. This energy budget seems to give a theoretical upper limit of energy release of a pulsar-like compact star.
@@ Acknowledgements.
The authors thank the anonymous referee for valuable suggestions. We are grateful to Shunshun Cao, Mingyu Ge, Kejia Lee, Bing Li, Ze-Nan Liu, Yudong Luo, Chenhui Niu, Yuan-Pei Yang, Xiaoping Zheng, Xia Zhou, and members of “DJ Find Girlfriend Assistant” Wechat group for helpful discussion. This work is supported by the National SKA Program of China (No. 2020SKA0120100) and the Strategic Priority Research Program of the CAS (No. XDB0550300). J.F.L. acknowledges support from the NSFC (Nos.11988101 and 11933004) and from the New Cornerstone Science Foundation through the New Cornerstone Investigator Program and the XPLORER PRIZE.